OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = x +

OpenStudy (nerdgirl):

I guess it'd just be 4 = 2x + 3. either that or switch the 4 and the 3 around.

OpenStudy (anonymous):

you cant replace y.....Y stays the same

OpenStudy (anonymous):

@StudyGurl14 @SolomonZelman

OpenStudy (studygurl14):

y=mx + b m = slope

OpenStudy (studygurl14):

then plug in point (3,4) to find b

OpenStudy (anonymous):

so it would be y=2x+ ? @StudyGurl14

OpenStudy (studygurl14):

right. but you have to find b to finish teh equation

OpenStudy (studygurl14):

y = 2x + b (3,4) (4) = 2(3) + b Solve for b, then plug back into y= 2x + b for teh final equation